Status Report

NASA Space Station On-Orbit Status 20 November 2004

By SpaceRef Editor
November 22, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 20 November 2004

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.  Saturday, the first of two off-duty days for the crew (except for standard maintenance and optional task list jobs).

Floating through their voluminous residence after wake-up at the regular 1:00am EST, morning inspection, hygiene and breakfast, the two crewmembers performed the regular weekly 3-hr. housecleaning.  [The “uborka stantsii” focuses on removal of food waste products, cleaning of compartments with vacuum cleaner, wet cleaning of the Service Module (SM) dining table and other surfaces with “Fungistat” disinfectant and cleaning of fan screens to avoid temperature rises.]

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

For today’s Saturday Science session CDR/SO Chiao reviewed the ISSI (In-Space Soldering Investigation) experiment and ran through the ISSI OBT (onboard training) and crew procedures for tests 4 and 5.  During a subsequent teleconference with the Principal Investigator (PI) via S-Band/audio, he discussed the experiment, to be conducted in the MWA (Maintenance Work Area) in the Lab.

At 9:00 am, the crew held the WPC (weekly planning conference) with the ground, discussing next week’s “Look-Ahead Plan” (prepared jointly by MCC-H and MCC-M planners), via S-band/audio.

Salizhan completed the regular daily maintenance/inspection of the Service Module (SM)’s environment control & life support systems (SOZh), including routine toilet system (ASU) servicing.

The crew performed their regular daily 2.5-hr. physical exercise program on TVIS treadmill, RED exerciser, and VELO cycle with bungee cord load trainer.

Afterwards, Chiao attended to the daily transfer of TVIS and RED exercise data files to the MEC (medical equipment computer) and also conducted the periodic transfer of accumulated data files from the wristband HRM (heart rate monitor) receiver stations to the MEC for downlink, followed by their erasure on the HRM.

At ~1:50pm, the CDR had his weekly PFC (private family conference) via S-band/audio and Ku-band/NetMeeting video.

Update on EVA battery reconditioning:  The EVA battery maintenance/reconditioning in the Airlock is continuing after its successful kickoff yesterday, and CDR Chiao had time set aside today for a checking up on the process.   [So far, the ground has seen expected signatures on all the batteries except for the EHIP (EMU helmet interchangeable portable light) battery installed in battery charger #2 (BC2), Channel 2, which looks “like it is either not installed, not fully seated, or somehow incompatible with its BC2 channel”.  Troubleshooting instructions were uplinked overnight for Leroy to check out these alternatives.]

Update on RPC-4 trip:  Yesterday’s checkout of the Lab UOP-4 (utility outlet panel #4) controlled by RPC-4 (remote power controller #4) showed a nominal current draw.  However, there is still a chance of an intermittent internal short in the UOP, and it will be left unused for potentially additional troubleshooting in the future.   [UOP-4 was the power outlet for the CEVIS (cycle ergometer with vibration isolation).  If the crew is able to clear the task-listed CEVIS continuity checks from the “job jar” today and they are good, they will get the Go to plug the cycle into Lab UOP-5 instead.]

Update on PL-2 MDM:  PL-2 MDM (payload multiplexer/demultiplexer #2) is back up and functioning nominally, now configured as unpowered backup.   [After a failure indication and loss of payload Ku-band telemetry on 11/13, analysis of dumped data showed corrupted control registers in a chip.  A power cycling of the PL-2 computer reinitialized the registers and cleared the failure indications.  The “taxi” chip in question on the PL-2’s “Hurdle” card (HRDL, high rate data link) converts electrical signals to optical signals for transmission over the fiber optic network to the Ku-band system for downlink.]

On Monday (11/22), a software patch will be uploaded to the GNC (guidance, navigation & control) MDMs to correct an error in the U.S. navigation software.   [To perform the load, the backup GNC2 will be reinitialized and recovered, leaving only the primary GNC1 in operation.  Once the backup MDM has been loaded and configured, attitude control will be handed over to the Russian Segment (RS).  The GNC MDMs will then be swapped and the new backup will be loaded with the patch, while the new primary will be reconfigured to regain attitude control from the RS.  The crew will be provided with a contingency procedure to help restore communications with MCC-H in the event that an unexpected primary GNC failure occurs during the patch load when there is no back-up GNC MDM.]

Weekly Science Update (Expedition Ten — 4th):

GASMAP:  Nothing new.

Human Research Facility/Workstation (HRF WS):  Continuing.

Advanced Ultrasound (ADUM):  The crew was thanked for last week’s highly successful scanning session.

Hand Posture Analyzer (HPA:  Nothing new.

In-Space Soldering Investigation (ISSI):  The ground was looking forward to today’s Saturday Science session with Science Officer Chiao.

Pore Formation and Mobility Investigation (PFMI):  Nothing new.

Space Acceleration Measurement System (SAMS):  SAMS continues collecting numerous channels of acceleration data that serve toward characterization of the dynamic vibratory environment on the ISS.

Microgravity Acceleration Measurement System (MAMS):  MAMS monitoring of the low-frequency acceleration environment of the space station continues.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):  “STES is in fine shape”.

Protein Crystal Growth Monitoring by Digital Holographic Microscope (PromISS):  Nothing new.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE):    Planned.

Binary Colloidal Alloy Test-3 (BCAT-3):  Nothing new..

Renal Stone (RS):  Nothing new.

Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES):  Nothing new.

Foot/Ground Reaction Forces during Space Flight (FOOT):  Nothing new.

Materials ISS Experiment (MISSE):  In progress. Deployed outside on the U.S. Airlock.  Nominal and collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI):  Nothing new.

Biotechnology Specimen Temperature Controller (BSTC):  In planning.

Yeast Group Activation Packs (Yeast GAP):  Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM):  Nothing new.

Earth Science Toward Exploration Research (ESTER):  Nothing new.

Serial Network Flow Monitor (SNFM):  The three SNFM captures on 11/17 were successful in capturing nine hours of the SAMS Drawer 1 bootup process.  Although the completion of bootup afterwards was not captured, sufficient data exists to give researchers good insight into the traffic characteristics.

Fluid Merging Viscosity Measurement (FMVM):  Planned.

Viscous Liquid Foam–Bulk Metallic Glass (Foam):   Nothing new.

Effects of Prolonged Space Flight on Human Skeletal Muscle (BIOPSY):  Nothing new.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2):  Planned.

Commercial Generic Bioprocessing Apparatus (CGBA):  Nothing new.

Miscible Fluids in Microgravity (MFMG):  Nothing new.

Educational Payload Operations (EPO):  Nothing new.

Capillary Flow Experiment (CFE):   Nothing new.

Crew Earth Observations (CEO):  The ground team has received and reviewed a total of 1,539 of the current crew’s CEO images through 11/18.  An unexpected, excellent image of Lake Valencia, Venezuela will be published on NASA/GSFC’s Earth Observatory website this weekend.  Careful study and enhancement of the image reveals interesting swirls of algal blooms in the lake’s waters, the result of pollutants in runoff from nearby cities and agricultural activities.  In response to crew inquiry: the ground has noted a “real improvement in your technique with regards to continuity and overlap of your photos.  Thanks for your extra effort here.  This will become particularly critical for sessions with the long lenses where the field of view is small and locating small features is most challenging”.  The crew was also thanked for a nice job on their images of the Patagonian Glaciers target region.  Despite poor weather conditions they managed some useful, between-the-clouds views of some rarely photographed glaciers there.

Today’s optional CEO photo targets, no longer limited in the current LVLH attitude, were Salamat Basin fans, Chad (weather was predicted to clear over the Salamat megafans for high resolution photography.  Mapping swaths along divergent drainages and wetland boundaries to the right of track will be useful for mapping individual fans), and Lake Nasser, Toshka Lakes, Egypt (this overpass provided an opportunity for high-resolution photography of these recently-filled lakes.  Mapping swaths along the shorelines are useful for assessing ecologic and hydrologic changes around the lakes).

CEO images can be viewed at these websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 10 crew visit:

Expedition 10 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Upcoming Key Events: 

  • Soyuz hot-fire thrusters test — 11/24;
  • Soyuz relocation DC1-FGB — 11/29;
  • Progress 15P undocking & destructive reentry — 12/22;
  • Progress 16P launch — 12/23;
  • Progress 16B docking — 12/25 (GMT);
  • EVA-12 — January 05 (1/25?).

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit  (as of this morning, 10:08am EST [= epoch]):

  • Mean altitude — 358.0 km
  • Apogee height — 361.0 km
  • Perigee height — 354.9 km
  • Period — 91.70 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0004496
  • Solar Beta Angle — -3.2 deg (magnitude increasing)
  • Orbits per 24-hr. day — 15.70
  • Mean altitude loss in last 24 hours — 128 m
  • Revolutions since FGB/Zarya launch (Nov. 98) — 34289


ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.